### v1

# 3D Anderson localization of light in disordered systems of dielectric particles

Abstract
We present the results of full wave numerical simulations of light transmission by layers of irregular dielectric particles that demonstrate 3D Anderson localization of light in disordered uncorrelated discrete media. Our simulations show that a high degree of disorder in a dense layer prevents transverse spreading of a propagating beam and allows anomalous conservation of its initial polarization state. A transition from the purely diffusive regime to a non-exponential time dependence when reaching the Ioffe-Regel criterion is observed in short pulse transmission measurements. The effect depends on the turbidity of the layer: increasing the volume fraction and the refractive index leads to the enhancement of the non-exponential behaviour caused by the disorder in the medium which is a clear signature of Anderson localization.
keywords:
disordered medium, light localization, light scattering
1 Introduction
Complete understanding and control of light propagation in disordered media can be advantageous in different applications in photonics like imaging and focusing in random media, random lasing, radiative transfer in particulate media and optical remote sensing [1, 2, 3, 4, 5]. However, this goal has not been achieved yet, in particular, for dense turbid media, due to the complexity of the problem for both analytical analysis and numerical simulations.
Light propagation in sparse media, in which scatterers are separated by distances larger than the wavelength, can be understood as a sequence of scattering events that are determined by the optical properties of individual particles, like single-scattering phase function and single-scattering albedo. If the number of such events is large enough, the optical transport process can be modelled using analytical multiple-scattering theories and the classic diffusion equation (e.g., [6, 7, 8]). In dense media with high concentration of scatterers, multiple scattering includes near-field interactions, and more complex transport phenomena may take place. For example, interference of counter-propagating waves may result in a weak localization and a coherent backscattering enhancement [9, 10]. Strong Anderson localization (AL) has been predicted for media with strong disorder and with multiple scattering not being suppressed by absorption [11, 12]. An important condition for the AL emergence is the sub-wavelength scale of coherent multiple scattering that can be expressed through the Ioffe-Regel criterion $kl^{*}<1$ for the mean free path length $l^{*}$ (where $k$ is the wavenumber) [13, 14]. This suggests a necessity of dense packing of particles in a disordered system in order to obtain AL. With high enough scattering strength of dielectric scatterers, that can be controlled by the variation of their refractive index, one can expect signatures of AL in the observations of light transmission by layers of particles. However, even non-absorbing layers with required properties and thickness $kL\gg 1$ can be quite opaque as it was already recognized by Anderson [11], which complicates the study of their transmission properties even further.
Generally, AL is characterized by a transition from diffusive propagation to a reduction and even complete halt of diffusion. An explanation involving trajectories with closed loops and increased return probability is often suggested [15, 16, 17]. However, the actual mechanism behind such dynamics has not been described yet in detail. The phenomenon is general in nature and can be observed for different kinds of waves, e.g., in acoustics [18], quantum systems [19, 20] and 2D photonic structures [21]. At the same time, the existence of 3D AL for electromagnetic waves in disordered uncorrelated media remains under debate. Experimental measurements of light transmission by layers of highly scattering particulate materials could shed light on the problem [17, 22, 23]. However, certain problems with precise control of the sample properties and accounting for the effects of absorption [24] and inelastic scattering [25, 26, 27] make this difficult. The most recent experimental results reported in [23] are debated because of ambiguous interpretation of the detected signature of AL that can be explained by fluorescence [27].
In theoretical studies of AL, one has to properly describe the disordered structure of a model medium and carefully take into account the multiple scattering near-field effects, including the polarization state of the scattered near field [28, 29, 16, 30]. Conventional multiple scattering theories cannot be applied here because of the small mean free path values at dense packing where the far-field single scattering properties of particles do not play role. In existing analytical models of the AL phenomenon, simplified descriptions of the media are used, e.g., a cloud of point scatterers [31] or a dielectric constant fluctuation around the average value $\epsilon(\vec{r})=\bar{\epsilon}+\delta\epsilon(\vec{r})$ [15]. If applied to the real ’white paint’-like 3D media, this can be insufficient.
Full wave numerical simulations can avoid approximations and unnecessary scattering effects considering pure interaction of the electromagnetic field with the target structure that may have arbitrary geometry. Modern numerical methods allow generation of particulate layer samples with high degrees of disorder, arbitrary shape of scatterers and packing densities [28, 32, 33]. At the same time, such an electromagnetic problem is multi-scale by definition. If the target object is much larger than the wavelength and has sub-wavelength structural features, accurate solutions require large computational meshes. This is a technical problem that can be resolved with the application of high-performance computing (HPC). However, only a few works on 3D numerical simulations and light localization in disordered and partially disordered structures have been reported [34, 35, 36, 37].
Observable characteristics of AL are the reduction of transverse spreading of the propagating light beam [38, 36, 34] and, in particular, time-delayed transmitted energy decay [39, 34]. These anomalies as unambiguous signatures of AL were found for the first time in a recent full wave numerical study of light transmission through 3D layers of densely packed perfect electric conductor (PEC) spheres [34]. That work suggests a possibility of the AL existence for electromagnetic waves in 3D in general. However, the case of dielectric non-absorbing spheres in layers with similar parameters and refractive indices up to $n=10$ showed only diffuse transport. Thus, the question whether AL can be observed for electromagnetic waves for disordered dielectric 3D structures remains an open problem to date.
In this work, we apply numerical simulations and HPC to study if conditions for AL can be created at $kl^{*}\lesssim 1$, according to the Ioffe-Regel criterion, and if AL signatures can be observed in light transmission by layers of dielectric particles. Reaching high enough degree of disorder at dense packing of irregular constituents, we observe a transition of classical diffusion characterised by the exponential time dependence of transmission $T(t)$ to AL with the $T(t)$ decay rate deviating from exponent. This behaviour is accompanied by the reduction of transverse spreading observed in the propagation of a focused beam. Thus, we demonstrate a possibility of AL in the ’white-paint’-like 3D systems of dielectric particles.
2 Numerical model
We apply the discontinuous Galerkin time-domain method (DGTD) [40, 32, 41, 42] to solve Maxwell’s equations for multi-particle layers illuminated by a plane wave as well as by a focused beam. For a qualitative benchmark of the DGTD light transmission computations for simplified structures we utilized the CST Microwave Studio software package, which is based on the Time Domain Finite Integration Technique (FIT) (see Supplementary material A.2).
We note that the $kl^{*}\lesssim 1$ criterion unconditionally requires dense packing of scatterers. Maximally random jammed packings of regular constituents, like spheres and regular polyhedra, always form hyper-uniform structures regardless of the packing algorithm [43, 44]. This kind of discrete media packed near their corresponding maxima, that depend on the type of constituents, have a reduced degree of disorder which influences their transmission properties (e.g., [45, 28] ). The problem can be solved by considering scatterers with random irregular shapes. To our knowledge, numerical simulations studying light localization in such systems have not been reported yet. Observation of AL in 3D layers of PEC spheres and its absence in the dielectric equivalent structures [34] provides us with an additional motivation to consider the case of irregular constituents.
Here, we simulate light transmission by mono-disperse layers with up to 20000 particles and use Gaussian random field shapes as constituents [28, 32, 46]. The Bullet physics engine [47] is applied to numerically pack them in systems with volume fractions up to $\rho=0.5$.
The constituent particle size is determined by the goal of satisfying the $kl^{*}\lesssim 1$ criterion and, possibly, to create conditions for light localization. The scale of free-space voids in a mono-disperse system becomes comparable to the particle size at large volume fractions. We take the particle size $kr=1.1$, where $r$ is the radius of the circumscribing sphere. In densely packed samples of compact irregular shapes their circumscribing spheres overlap making the distances between neighbour particles sufficiently small.
We consider two types of cylindric layers: thick ones for studying the focused beam propagation and thinner, but horizontally more extended ones for time-resolved transmission measurements. In the first case, the diameter and thickness are $kR=22.1$ and $kL=19.5$, respectively, and the number of particles is 20000 (Fig. 1a). In the second case, layers with various packing densities $\rho$ have larger diameter to thickness ratios and numbers of particles up to 19000. A layer sample with $kR=32$, $kL=14.5$ and $\rho=0.44$ and its vertical cross-section are shown in (Figs. 3ab). We note that the sample sizes are limited by our computational resources and the complexity of the unstructured mesh generation. For the DGTD simulations, we generate meshes with the particle systems surrounded by the perfectly matching layer (PML) to simulate an open boundary. We do not apply periodic boundary conditions in order to avoid possible uncertainties caused by the periodicity of the mesh. Cylindrical symmetry of samples may also affect the result. Therefore, we use samples with two types of the side boundaries: a randomly perturbed one and a smooth boundary. In the first case defects of the scale of $\sim\lambda$ that break such symmetry are introduced. In the second case a smooth boundary forms a dielectric discontinuity between the disordered medium and free space and, generally, one can expect whispering-gallery (WG) modes due to total internal reflection and cylindrical symmetry [48].
We consider a non-absorbing dielectric material and vary the refractive index $n$ in the range of $2.0\leq n\leq 3.0$.
3 Results and discussion
Figure 1: Propagation of a continuous, focused beam through a dense and a sparse layers of irregular particles. (a) and (c), samples with volume fractions of 0.48 (dense) and 0.16 (sparse), respectively. (b) and (d), steady-state near-field intensity distributions $|E|^{2}$ in XY, YZ and back-XZ planes for both layers. A white dotted line indicates a layer boundary with the diameter $kR=22.1$ and the thickness $kL=19.5$. The incident beam is $E_{x}$-polarized. The size of particles is $kr=1.1$, and the refractive index is $n=3.0$. The sparse structure shows a classic multiple-scattering pattern while dense packing strongly limits transverse spreading of the beam. An animation of a short pulse propagation, illustrating the transport mechanism in the dense case, is provided in the Supplementary material A.1.
Diffusive propagation of light is characterized by a transverse spread of energy and exponential energy decay in time as a result of multiple scattering. The geometry of the field spread experiences a change when particles are densely packed. This effect depends on the particle size and refractive index. Generally, at dense packing, far-field single scattering becomes impossible, and classic multiple scattering turns into percolation of light along random transmission channels [28, 29, 30]. Depending on the particle size and degree of disorder, this can lead to formation of a Caylee-tree-like channel [28] or a transversely localized propagation [38, 36]. This may occur even at a relatively small refractive index $n=1.5$ and $l^{*}\lesssim\lambda$ [28], which, however, should be insufficient for AL. Nevertheless, the diffuse field component appears to be significantly reduced under such conditions, which can be seen in the polarimetric measurements of backscattering from high-albedo powders with different packing densities [49, 50, 51]. Little free space between the neighboring scatterers in a medium reduces the role of diffraction after each scattering event and, consequently, the contribution of the diffuse field [28]. The measured polarimetric features produced by single particles on the surface, that are suppressed by unpolarized diffuse multiple scattering in sparse samples, get enhanced in compressed ones. Hence, in the search for real AL, we consider samples with parameters that are close to those predicted in the theoretical works.
In Fig. 1, we study focused beam propagation through a thick layer of particles with a size $kr=1.1$ and a refractive index $n=3.0$. The internal field patterns are compared for a dense and a sparse structures of $\rho=0.48$ (Fig. 1a) and 0.16 (Fig. 1c), respectively. A continuous $E_{x}$-polarized focused beam incident on the front side in each case propagates in $Z$ direction. Steady-state intensity distributions as cross-sections in different planes are shown in Figs. 1b and 1d. The ”XY” panels correspond to the cross-sections near the back sides of the samples.
In the sparse case (Fig. 1d), we observe classic scattering by particle groups and individual particles, which appear to be weak scatterers at large separations from each other. In a dense system, the size of voids is reduced to $kl^{*}\lesssim 1$, and a focused beam tends to preserve its diameter along the path, re-emitting transmitted energy in the center of the back side. Small isotropic spreading occurs due to the leakage of energy through the evanescent field coupling. Therefore, the intensity of the beam decreases, but its geometric cross-section remains constant at the distance of the layer thickness. The dynamics of the process can be seen in the animation in the Supplementary material A.1. It shows the evolution of a focused beam, percolating through a dense layer without transverse spreading. Light propagation evolves mainly through evanescent coupling between individual scatterers without conventional free-space transport of the scattered waves. This reminds us of an evolution of a cellular automata (CA) in which the vector field state in an elementary cell at each time step is determined by its nearest environment, i.e. the shape of neighbour particles or the shape of a sub-wavelength void that is formed by the neighbors. The laws of interaction of the vector field and the air-dielectric interface become the CA rules. Then, a sequence of such interactions determines the evolution of the entire vector field from the initial state, resulting in a forward pulse propagation.
The fundamental property of the observed transport mechanism is that most of the energy of the propagating beam is concentrated in the free-space cavities, and little is refracted into the dielectric particles. A wave can couple to the neighbor-particle interface while being still coupled to the previous one propagating in a void between two interfaces. On the other hand, these interfaces are formally disconnected and belong to different scatterers. Thus, the field appears to be constantly coupled and one can say that there are no more ”scattering events”. Therefore, the term ”free path length” becomes irrelevant in its classic definition (see also [14]). $kl^{*}<1$ implies also an extremely small mean free time $t^{*}$, and the near field, being coupled to a particle, most often reaches its neighbour at time $t\ll T$, where $T$ is the wave oscillation period. Thus, we refer to the $kl^{*}$ parameter only to characterise the size of the voids in a particulate medium with respect to the wavelength.
Furthermore, we note that the wave packet is partially preserved along the propagation path. Its effective wavelength becomes smaller, $\lambda_{eff}\approx 0.6\lambda_{inc}$, increasing the effective size of particles to $k_{eff}r\approx 1.8$. Nevertheless, the condition $kl^{*}\lesssim 1$ is preserved for most of the voids.
Figure 2: Propagation of a continuous pulse through the dense layer shown in Fig. 1a. Steady state distribution of the components of the electric near field $\vec{E}$ (XZ-plane) show that the transmitted pulse is not totally depolarized but, on the contrary, preserves most of its initial polarization state $E_{x}$ after transmission. The $E_{y}$ component as well as $E_{z}$ is excited locally in the voids between the particles along the propagation path but is not transmitted directionally.
Diffusive propagation in a disordered medium depolarizes incident radiation after many scattering events that are strongly stochastic. A steady-state distribution of the electric field component $E_{x}$ in Fig. 2 computed for a dense layer shows a counter-intuitive conservation of the original polarization state of the propagating beam at the exit. The $E_{y}$ and $E_{z}$ components are excited along the path and play a role in the field coupling process but they do not propagate directionally. This fact can explain the halt of transverse spreading: the propagating beam ”prefers” cavities where the initial polarization direction can be preserved and, hence, the $\vec{k}$-vector bending is avoided and the direction of propagation is not changed.
Figure 3: (a) A layer of 19000 irregular particles packed with a volume fraction of 0.44 and with dimensions of $kR=32$ and $kL=14.5$. The size of particles is $kr=1.1$ and the refractive index is $n=3.0$. (b) vertical cross-section of the layer, showing intentional defects on the side boundary and the characteristic scale of voids $kl\lesssim 1$. (c) Normalized transmission of a short pulse by layers with volume fractions from $\rho=0.17$ to $0.5$ as a function of time $T(t)$, exhibiting a transition from the exponential regime to a non-exponential decay at longer times. (d) Time-dependent diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$. Both axes have logarithmic scale. A dense layer with $\rho=0.44$ demonstrates a possible transition to $t^{-1}$ (dashed line) dependence at times $t>20t_{D}$.
Figure 4: Simulation results for layers with different volume fractions and smooth side boundaries. The sample dimensions are of $kR=26$ and $kL=12$, respectively. The size of particles is $kr=1$ and the refractive index is $n=3.0$. (a) Normalized transmission of a short pulse by layers with volume fractions from $\rho=0.1$ to 0.5 as a function of time $T(t)$, exhibiting a transition from the purely exponential regime to a non-exponential decay at $\rho\gtrsim 0.3$. (b) Time-dependent diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$. Dense layers demonstrate a $t^{-1}$ (dashed lines) dependence at times $t>20t_{D}$. (c) and (d), Snapshots of the internal electric field (XY-cross-sections, $E_{x}$ component) distributions showing a whispering-gallery wave propagating along the boundary at time $t=25t_{D}$ for a dense ($\rho=0.44$) and a moderately sparse ($\rho=0.27$) layers. In the last case $T(t)$ is exponential despite the whispering-gallery effect.
Transverse localization in a dense structure may be accompanied by the changes in the time dependence of transmission, analogous to those observed for dense packings of PEC spheres in [34]. Here we perform a related analysis. We simulate time-resolved transmission by layers $T(t)$ varying their turbidity, i.e. the volume fraction and the refractive index. At first, we consider layers with perturbed side boundaries and with dimensions of $kR=32$ and $kL=14.5$. The volume fraction $\rho$ is varied between 0.17 and 0.5.
[…]
Non-exponential decay can be formally represented by means of a time-dependent diffusion coefficient $D(t)$. At the localized state, it decreases as $t^{-1}$ according to the solution of the diffusion equation for the case of $kl^{*}<1$ in time domain [39]. Figure 3d shows time resolved diffusion coefficients $D(t)$ deduced from the computed $T(t)$ by local exponential fitting. One can see a transition from small variation of $D(t)$ in sparse structures to a steep function at $\rho>0.44$ and times $t>15t_{D}$.
Evidently we would have to extend the observation time limit in order to study the evolution of $T(t)$ at longer times and to find out if $D(t)$ can reach the predicted $t^{-1}$ dependence. This could be achieved by increasing sample thickness and diameter roughly by a factor of two or more. This would correspond to $kL\gtrsim 30$ and a number of particles $N>10^{5}$ which makes the problem computationally extremely expensive, even for modern large scale HPC clusters. Another way to hinder energy decay in time is to create conditions for a whispering-gallery effect at the side boundary of a cylindric layer. For this purpose we did computations for layers with smooth boundaries and a particle size of $kr=1$. The results for $T(t)$ and $D(t)$ dependencies at different volume fractions from $0.1$ to $0.5$ are shown in Figures 4ab. The dimensions of the layers are $kR=26$ and $kL=12$. A densely packed sample with $\rho=0.44$ is shown in the inset in Figure 4a. Figures 4cd represent snapshots of the electric field distributions ($E_{x}$ field component) in the horizontal XY-plane at time $t=25t_{D}$ for a dense ($\rho=0.44$) and a moderately sparse ($\rho=0.27$) layers. The WG waves, excited by the incident pulse due to the smooth boundaries of samples with cylindric symmetry, are coupled to the side boundary due to total internal reflection and allow us reaching $t\approx 40t_{D}$ for relatively small samples. At the same time the field distributions remain random in the bulk. Despite the WG effect, present in both cases, the structures show different transmission properties. The $T(t)$ plot in Figure 4a clearly shows a transition from purely exponential decay to localization at large volume fractions and times $t>20t_{D}$. This occurs at $\rho\gtrsim 0.3$. The WG effect is observed in all the layers above and below this value, even in the sparse structures for which the the condition $l^{*}\lesssim\lambda$ holds. This proves that it is not the WG mode itself but the particle packing density which is causing the non-exponential behaviour of $T(t)$. The diffusion coefficient $D(t)$ at times $t\gtrsim 20t_{D}$ obtained for the dense structures with $\rho\gtrsim 0.44$ follows a $t^{-1}$ fit in Figure 4b demonstrating that we have reached the AL regime. We note that this result is consistent, in terms of time and the volume fraction threshold, with that obtained for layers of PEC spheres in [34].
Figure 5: (a) Normalized transmission of a short pulse through a layer with a volume fraction $\rho=0.44$ and a smooth boundary as a function of time $T(t)$ at different refractive indices. (b) Time-dependent diffusion coefficient $D(t)$ as obtained by a local exponential fitting of $T(t)$. A localization signature $D(t)\sim t^{-1}$ emerges at times $t>20t_{D}$ and $n=3.0$.
[…]
Figure 6: A snapshot of the near field $|E|^{2}$ excited by the propagation of a short pulse through a dense layer with $\rho=0.48$ and $n=3.0$ shown in Fig. 1a along $Z$ axis at times $t=2t_{D}$ (a) and $t=20t_{D}$ (b). A fraction of light is trapped by the disordered air-dielectric interface through evanescent coupling. This results in a constant flow along complex, highly disordered paths which slows down the energy decay through the layer boundaries. An animation showing the time dynamics of the near field can be found in the Supplementary material A.1.
Our results imply a mechanism that slows down the energy decay at dense packing. Apparently, it should be connected with the percolation-like light transport in a thick, dense layer we analyzed above. Visualization of the field dynamics excited by a short pulse yields a clue to this mechanism. One can conventionally split up the propagating field into two components. The first one is the wave packet of the incident pulse that quickly diffuses through a layer. The second one is created by the energy leaked from the packet through the evanescent field-coupling. Being coupled once, this component remains trapped by the air-dielectric interface, which, together with the free-space cavities, represents a geometrically complex 3D structure. Despite little free space at dense packing, such a structure possesses a long-range connectivity, i.e. arbitrary points in a void can be connected with another point in an arbitrary free-space location. The paths connecting such arbitrary points can be much longer than the sample dimensions.
In Fig. 6, one can see snapshots of the near-field intensities inside the thick layer shown in Fig. 1a at a time right after the pulse passage $t=2t_{D}$ and at the longer time $t=20t_{D}$. An animated field evolution for a period of time after $2t_{D}$ can be found in the Supplementary material A.1. The trapped field uniformly fills the sample volume and propagates along complex paths, constantly changing direction and forming hotspots in the voids. The paths can be not necessarily closed loops, but they are apparently sufficiently long. The field decay occurs only when the field reaches the sample boundaries. Thus, the energy in the interior needs substantial time in order to reach the boundary, much longer than the time of propagation along a straight line or a sequence of straight lines between scattering events in a sparse medium. The character of such an internal field evolution does not significantly change at longer times, and the decay of hotspots is not monotonous in each location. They demonstrate different behaviour in such a disordered system: those that are excited directly by the transmitted pulse decay while others arise in new locations. The decay rate is also different for different hotspots, i.e. they may have different life times. Some are decaying within $\approx 5$ oscillation periods while others are much more stable. Non-monotonous decay and enhancement with time may also take place in the same location, which proves a high geometrical complexity of the paths. If the energy of a path is leaked at the sample boundary, it is lost in the total decay. A leakage in the interior transfers energy to new paths with similar spatial scales and degrees of complexity, forming a complex percolation graph. Thus, the entire process appears to be a sophisticated spatiotemporal 3D flow of energy and should be a subject for further studies. In particular, such a mechanism can be relevant for low-absorbing materials as it involves field-interface interaction and propagation in free space. The role of the refractive index also becomes clear in this situation. It determines the balance between the energy transmitted by particles and chains of particles, enhancing a diffusive process, and the energy flowing through a random system of connected cavities with long path lengths that participates in localization.
4 Conclusions
We observe clear signatures of AL in the results of the full wave numerical simulations of light transmission by 3D dielectric particulate layers with high degrees of structural disorder. A transition from diffusive transport to localization occurs at approaching the Ioffe-Regel criterion for the mean free path length $kl^{*}\lesssim 1$ in densely packed systems and at the refractive index of $n=3.0$, which is consistent with the predictions of theoretical models. Simulation of the focused beam propagation in a dense layer demonstrates an absence of transverse spreading and anomalous conservation of the beam’s linear polarization state in contrast to the classical light scattering in a sparse analog structure. Time-resolved transmission of a short pulse in layers with similar densities and constituent particle sizes exhibits a non-exponential energy decay already at times $t>15t_{D}$. For observations at $t\gg t_{D}$ larger structures must be considered which makes the problem computationally very expensive even using HPC. Alternatively, we extend the observation time retaining sample cylindric symmetry. This allows excitation of whispering gallery modes that hinder field decay from a sample. Therewith we found that the time-resolved diffusion coefficient varies in agreement with the self-consistent theory of localization at times $t>20t_{D}$ and volume fractions $\rho\gtrsim 0.44$. Thus, Anderson localization of light is possible in dense highly disordered 3D dielectric structures which motivates a continued experimental search of the AL characteristics in uncorrelated disordered media, with realistic material refractive indices not exceeding $n=3.0$.
[…]
Methods We performed numerical simulations using a self-developed light scattering code based on the open source DGTD Maxwell solver MIDG published by Tim Warburton [52]. The CST Microwave Studio software package based on the Time Domain Finite Integration Technique was used for a qualitative benchmarking the DGTD simulation results for simplified setups. In our code we implemented plane-wave and focused beam lights sources, boundary conditions and material description. Here, we use continuous and Gaussian pulses. A plane wave source is realized using the total field/scattered field technique. For simulations with a focused beam we implemented the plane wave superposition method [53]. An open computational domain is pretended with the PML absorbing boundary condition. For the time-resolved transmission measurements we illuminated target layers with a linearly polarized short plane wave pulse. The intensity of the transmitted signal is recorded, correspondingly, in a monitor plane behind the layer. The intensity curves in all plots are normalized by the peak intensity of the transmitted pulse. For generation of the particulate samples with different dimensions and controlled filling fractions we use the Bullet Physics engine [47]. This is an open-source C++ library that can be also used as an add-on in the Blender software. It allows simulation of dynamics and collision of multiple arbitrary 3D shapes in time domain. Thus, simulating natural powder mechanics samples with realistic topologies can be created. For sparse distributions of the pre-generated random irregular shapes we simulated their free fall on a substrate in a closed cylindrical volume. The sets of particles are created with a self-developed shape generator based on the Gaussian random field approach [46]. 3D models of the generated layers were then used for tetrahedral mesh generation. The spatial resolution in the regions with dielectric material with $n=3.0$ was $\sim 15$ tetrahedral cells per central wavelength of the incident pulse and the nodal expansion order in the DGTD numerical scheme was $N=3$. All the DGTD simulations were done on the HPC cluster Noctua 2 of Paderborn University using up to 30 cluster nodes per simulation.
Appendix A Supplementary material
A.1 Near field dynamics
Animated evolution of the near field in time helps to better represent the field dynamics in such a complex structure as a discrete random medium consisting of densely packed irregular particles. To support our analysis of the static field snapshots in Figure 1b we additionally simulate propagation of a short focused pulse in the same dense layer sample (Figure 1a) with the same model parameters. Animation A.1 represents an $XZ$ cross-section of the computational domain with the near-field intensity distribution and shows a cellular-automata-like behavior of the local field hotspots in the sub-wavelength voids between the closely neighboring particles.
Figure A.1: Animated propagation of a short focused pulse in a dense layer shown in Figure 1a.
Animation A.2 demonstrates dynamics of the near field that is trapped by the air-dielectric interface of the constituent particles after propagation of a short plane-wave pulse in a short period near $t=2t_{D}$. At high degree of disorder and in the absence of large voids the evanescent coupling mechanism slows down the energy decay from the system.
Figure A.2: Animated dynamics of the near field in a dense layer shown in Figure 1a excited by a short plane wave pulse.
A.2 Pulse transmission by layer samples with reduced dimensions
Here we simulate short-pulse transmission by layers with reduced dimensions in order to study the transition to AL increasing the geometrical complexity through the layer thickness and the number of particles. In addition we do test simulations using the CST Microwave Studio software package. It is based on the Time Domain Finite Integration Technique, so, one can qualitatively benchmark our DGTD results with a different numerical method and a different electromagnetic solver.
We take a cylindrically symmetric sample with $kR=26$, $kL=12$ and $\rho=0.44$ as a reference and consider reduced structures with thicknesses $kL=6$ and 3 and an elementary bi-layer structure. The refractive index is fixed at $n=3.0$. The results of the simulated transmission $T(t)$ for four layers are presented in Figure A.3a. A bi-layer shows purely exponential decay even though a WG mode is excited. The case of $kL=3$ is equivalent to a doubling the thickness of the bi-layer sample. This appears to be enough to make the $T(t)$ curve non-exponential. In all four cases a WG mode is excited allowing us reaching the $25-40t_{D}$ region even with such a small thickness. Further thickness increasing leads to stronger deviations of exponent and faster reduction of the diffusion coefficient $D(t)$ (Figure A.3b). However, only with $kL=12$ we obtain a qualitative change at longer times and reach the $t^{-1}$ fit.
Figure A.3: (a) Normalized transmission of a short pulse through layers with cylindric symmetry and different thickness $kL$ as a function of time $T(t)$. The diameter and volume fraction are $kR=26$ and $\rho=0.44$, correspondingly, and the refractive index is $n=3.0$. The energy decay from a simple double layer of particles is exponential while thicker layers demonstrate increasingly non-exponential transmission. (b) Diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$. $D(t)$ quickly evolves from a constant to a $t^{-1}$ fit (dashed line) with increasing thickness. (c) Vertical cross-sections of the samples with different thickness.
Figure A.4: Normalized transmission of a short pulse through a small layer of 2000 particles packed with volume fraction $\rho=0.5$ (inset) and with dimensions $kR=15$ and $kL=3$ as a function of time $T(t)$ at different refractive indices. Simulations are done with the CST Microwave Studio software package.
For the computations with the CST FIT solver we take a cylindric sample with thickness $kL=3$ (a four-particle thick layer) and reduced diameter $kR=15$. Due to parallelization restrictions of CST MWS we need to simplify the target system. With the particle size $kr=1$ and the volume fraction $\rho=0.5$ the number of particles becomes relatively small in this case, N=2000. With fixed sample parameters we vary the material refractive index $n$ from 2.0 to 3.0. For the CST computations we use a polyhedral description of the model sample and, then, a hexahedral mesh is generated by the software internally. The electromagnetic problem is solved in time domain simulating propagation of the short plane-wave pulse, similarly to our DGTD method setup.
The result in Figure A.4 is qualitatively similar to the $T(t)$ refractive index dependence in Figure 5a obtained for a much larger system with the DGTD method.
Summarizing all our results we note that increasing three parameters: layer thickness, volume fraction and refractive index act in very similar ways causing transition from an exponential dependence of $T(t)$ to a non-exponential behavior with qualitative change at longer times.

### v4

@@
 Abstract
-We present the results of full wave numerical simulations of light transmission by layers of irregular dielectric particles that demonstrate 3D Anderson localization of light in disordered uncorrelated discrete media.
-Our simulations show that a high degree of disorder in a dense layer prevents transverse spreading of a propagating beam and allows anomalous conservation of its initial polarization state.
-A transition from the purely diffusive regime to a non-exponential time dependence when reaching the Ioffe-Regel criterion is observed in short pulse transmission measurements.
-The effect depends on the turbidity of the layer: increasing the volume fraction and the refractive index leads to the enhancement of the non-exponential behaviour caused by the disorder in the medium which is a clear signature of Anderson localization.
-keywords:
-disordered medium, light localization, light scattering
-1 Introduction
-Complete understanding and control of light propagation in disordered media can be advantageous in different applications in photonics like imaging and focusing in random media, random lasing, radiative transfer in particulate media and optical remote sensing [1, 2, 3, 4, 5].
-However, this goal has not been achieved yet, in particular, for dense turbid media, due to the complexity of the problem for both analytical analysis and numerical simulations.
-Light propagation in sparse media, in which scatterers are separated by distances larger than the wavelength, can be understood as a sequence of scattering events that are determined by the optical properties of individual particles, like single-scattering phase function and single-scattering albedo.
-If the number of such events is large enough, the optical transport process can be modelled using analytical multiple-scattering theories and the classic diffusion equation (e.g., [6, 7, 8]).
-In dense media with high concentration of scatterers, multiple scattering includes near-field interactions, and more complex transport phenomena may take place.
-For example, interference of counter-propagating waves may result in a weak localization and a coherent backscattering enhancement [9, 10].
-Strong Anderson localization (AL) has been predicted for media with strong disorder and with multiple scattering not being suppressed by absorption [11, 12].
-An important condition for the AL emergence is the sub-wavelength scale of coherent multiple scattering that can be expressed through the Ioffe-Regel criterion $kl^{*}<1$ for the mean free path length $l^{*}$ (where $k$ is the wavenumber) [13, 14].
-This suggests a necessity of dense packing of particles in a disordered system in order to obtain AL.
-With high enough scattering strength of dielectric scatterers, that can be controlled by the variation of their refractive index, one can expect signatures of AL in the observations of light transmission by layers of particles.
-However, even non-absorbing layers with required properties and thickness $kL\gg 1$ can be quite opaque as it was already recognized by Anderson [11], which complicates the study of their transmission properties even further.
-Generally, AL is characterized by a transition from diffusive propagation to a reduction and even complete halt of diffusion.
+We present the results of full-wave numerical simulations of light transmission through layers of irregular dielectric particles, demonstrating three-dimensional Anderson localization of light in disordered, uncorrelated discrete media.
+Our simulations show that a high degree of disorder in a dense layer suppresses the transverse spreading of a propagating beam.
+A transition from the purely diffusive regime to a non-exponential temporal dependence is observed in short-pulse time-resolved transmission measurements as the system approaches the Ioffe-Regel condition.
+Along with this, near-field dynamics leads over time to the formation of spatially localized modes and the transmission spectrum becomes consistent with the Thouless criterion.
+The effect depends on the turbidity of the layer: increasing the volume fraction of scatterers and the refractive index contrast enhances the non-exponential behavior induced by disorder, which is a clear signature of Anderson localization.
+preprint: APS/123-QED
+A complete understanding and control of light propagation in discrete disordered media can be advantageous in different applications in photonics, like imaging and focusing in random media, random lasing, radiative transfer and optical remote sensing [1, 2, 3, 4, 5].
+However, modeling light transport phenomena in highly disordered turbid media remains a challenge due to the inherent complexity and multi-scale nature of the problem.
+While classic wave diffusuion and weak localization, observed as coherent backscattering enhancement [6, 7, 8], are well understood, strong Anderson localization of light (AL) in 3D structures is a subject of ongoing research.
+AL has been predicted for media with strong disorder and strong multiple scattering [9, 10].
+Generally, it is characterized by a transition from diffusive propagation to a reduction and even complete halt of diffusion.
 An explanation involving trajectories with closed loops and increased return probability is often suggested [11, 12, 13].
 However, the actual mechanism behind such dynamics has not been described yet in detail.
-The phenomenon is general in nature and can be observed for different kinds of waves, e.g., in acoustics [18], quantum systems [19, 20] and 2D photonic structures [21].
-At the same time, the existence of 3D AL for electromagnetic waves in disordered uncorrelated media remains under debate.
+An important condition for the AL emergence is the sub-wavelength scale of coherent multiple scattering that can be expressed through the Ioffe-Regel criterion $kl^{*}\lesssim 1$ for the mean free path length $l^{*}$ (where $k$ is the wavenumber) [14, 15].
+The phenomenon is general in nature and can be observed for different kinds of waves, e.g., in acoustics [16], quantum systems [17, 18], 2D [19, 20] and spatially correlated 3D photonic structures [21].
+The existence of AL for electromagnetic waves in 3D uncorrelated media and metallic materials has been demonstrated recently through a numerical study [22].
+Observable characteristics of AL are the reduction of transverse spreading of the propagating light beam [23, 24, 22] and, in particular, time-delayed transmitted energy decay [25, 22].
+Transmission spectrum in this case is characterized by non-overlapping sharp peaks [26, 27, 28, 29].
+Numerical modeling of light transmission through layers of overlapping metallic spheres revealed these anomalies as unambiguous signatures of AL [22], suggesting its general possibility in 3D.
+At the same time analogous dielectric structures did not exhibit the phenomenon [22].
+Thus, the question whether AL can be observed for electromagnetic waves and disordered dielectric 3D media remains an open problem to date.
 Experimental measurements of light transmission by layers of highly scattering particulate materials could shed light on the problem [13, 30, 31].
 However, certain problems with precise control of the sample properties and accounting for the effects of absorption [32] and inelastic scattering [33, 34, 35] make this difficult.
-The most recent experimental results reported in [23] are debated because of ambiguous interpretation of the detected signature of AL that can be explained by fluorescence [27].
-In theoretical studies of AL, one has to properly describe the disordered structure of a model medium and carefully take into account the multiple scattering near-field effects, including the polarization state of the scattered near field [28, 29, 16, 30].
-Conventional multiple scattering theories cannot be applied here because of the small mean free path values at dense packing where the far-field single scattering properties of particles do not play role.
-In existing analytical models of the AL phenomenon, simplified descriptions of the media are used, e.g., a cloud of point scatterers [31] or a dielectric constant fluctuation around the average value $\epsilon(\vec{r})=\bar{\epsilon}+\delta\epsilon(\vec{r})$ [15].
-If applied to the real ’white paint’-like 3D media, this can be insufficient.
-Full wave numerical simulations can avoid approximations and unnecessary scattering effects considering pure interaction of the electromagnetic field with the target structure that may have arbitrary geometry.
-Modern numerical methods allow generation of particulate layer samples with high degrees of disorder, arbitrary shape of scatterers and packing densities [28, 32, 33].
-At the same time, such an electromagnetic problem is multi-scale by definition.
-If the target object is much larger than the wavelength and has sub-wavelength structural features, accurate solutions require large computational meshes.
-This is a technical problem that can be resolved with the application of high-performance computing (HPC).
-However, only a few works on 3D numerical simulations and light localization in disordered and partially disordered structures have been reported [34, 35, 36, 37].
-Observable characteristics of AL are the reduction of transverse spreading of the propagating light beam [38, 36, 34] and, in particular, time-delayed transmitted energy decay [39, 34].
-These anomalies as unambiguous signatures of AL were found for the first time in a recent full wave numerical study of light transmission through 3D layers of densely packed perfect electric conductor (PEC) spheres [34].
-That work suggests a possibility of the AL existence for electromagnetic waves in 3D in general.
-However, the case of dielectric non-absorbing spheres in layers with similar parameters and refractive indices up to $n=10$ showed only diffuse transport.
-Thus, the question whether AL can be observed for electromagnetic waves for disordered dielectric 3D structures remains an open problem to date.
-In this work, we apply numerical simulations and HPC to study if conditions for AL can be created at $kl^{*}\lesssim 1$, according to the Ioffe-Regel criterion, and if AL signatures can be observed in light transmission by layers of dielectric particles.
-Reaching high enough degree of disorder at dense packing of irregular constituents, we observe a transition of classical diffusion characterised by the exponential time dependence of transmission $T(t)$ to AL with the $T(t)$ decay rate deviating from exponent.
-This behaviour is accompanied by the reduction of transverse spreading observed in the propagation of a focused beam.
-Thus, we demonstrate a possibility of AL in the ’white-paint’-like 3D systems of dielectric particles.
-2 Numerical model
-We apply the discontinuous Galerkin time-domain method (DGTD) [40, 32, 41, 42] to solve Maxwell’s equations for multi-particle layers illuminated by a plane wave as well as by a focused beam.
-For a qualitative benchmark of the DGTD light transmission computations for simplified structures we utilized the CST Microwave Studio software package, which is based on the Time Domain Finite Integration Technique (FIT) (see Supplementary material A.2).
-We note that the $kl^{*}\lesssim 1$ criterion unconditionally requires dense packing of scatterers.
-Maximally random jammed packings of regular constituents, like spheres and regular polyhedra, always form hyper-uniform structures regardless of the packing algorithm [43, 44].
-This kind of discrete media packed near their corresponding maxima, that depend on the type of constituents, have a reduced degree of disorder which influences their transmission properties (e.g., [45, 28] ).
+In this work, we apply numerical simulations and high-performance computing to demonstrate that AL signatures can be observed in light transmission by 3D layers of irregular dielectric particles.
+Reaching high enough degree of spatial disorder, we obtain a transition of classical diffusion characterized by the exponential time dependence of transmission $T(t)$ to AL with the $T(t)$ decay rate deviating from exponent.
+This behavior is accompanied by the reduction of transverse spreading observed in the propagation of a focused beam.
+We note that the problem requires proper description of the disordered geometry of a model medium and carefull accounting for the multiple scattering near-field effects, including the polarization state of the scattered near field [36, 37, 12, 38, 39].
+In existing analytical solutions, simplified descriptions of the media are used (e.g., [39, 11, 40]) which can be insufficient if applied to the vector electromagnetic field in real ’white paint’-like 3D media.
+With full wave numerical simulations one can avoid approximations and consider pure electromagnetic scattering in the target structure with arbitrary geometry [36, 41, 42, 22, 43, 24, 21].
+We use the discontinuous Galerkin time-domain method (DGTD) [44, 41, 45, 46] to solve Maxwell’s equations for multi-particle layers illuminated by a plane wave as well as by a focused beam (see details in Supplementary material I).
+For a qualitative benchmark of the DGTD computations for simplified structures we utilized the CST Microwave Studio software package, which is based on the Time Domain Finite Integration Technique (FIT) (see Supplementary material II).
+The $kl^{*}\lesssim 1$ criterion unconditionally requires dense packing of scatterers.
+Maximally random jammed packings of regular constituents, like spheres and regular polyhedra, always form hyper-uniform structures regardless of the packing algorithm [47].
+This kind of discrete media packed near their corresponding maxima have a reduced degree of disorder which influences their transmission properties (e.g., [48, 36] ).
 The problem can be solved by considering scatterers with random irregular shapes.
-To our knowledge, numerical simulations studying light localization in such systems have not been reported yet.
-Observation of AL in 3D layers of PEC spheres and its absence in the dielectric equivalent structures [34] provides us with an additional motivation to consider the case of irregular constituents.
 Here, we simulate light transmission by mono-disperse layers with up to 20000 particles and use Gaussian random field shapes as constituents [36, 41, 49].
 The Bullet physics engine [50] is applied to numerically pack them in systems with volume fractions up to $\rho=0.5$.
-The constituent particle size is determined by the goal of satisfying the $kl^{*}\lesssim 1$ criterion and, possibly, to create conditions for light localization.
+The constituent particle size is determined by the goal of satisfying the $kl^{*}\lesssim 1$ criterion.
 The scale of free-space voids in a mono-disperse system becomes comparable to the particle size at large volume fractions.
-We take the particle size $kr=1.1$, where $r$ is the radius of the circumscribing sphere.
+We take the particle size $X_{r}=kr=1.1$, where $r$ is the radius of the circumscribing sphere.
 In densely packed samples of compact irregular shapes their circumscribing spheres overlap making the distances between neighbour particles sufficiently small.
 We consider two types of cylindric layers: thick ones for studying the focused beam propagation and thinner, but horizontally more extended ones for time-resolved transmission measurements.
-In the first case, the diameter and thickness are $kR=22.1$ and $kL=19.5$, respectively, and the number of particles is 20000 (Fig.
+In the first case, the dimensionless size parameters of the diameter and thickness are $X_{R}=22.1$ and $X_{L}=19.5$, respectively, and the number of particles is 20000 (Fig.
 1a).
-In the second case, layers with various packing densities $\rho$ have larger diameter to thickness ratios and numbers of particles up to 19000.
-A layer sample with $kR=32$, $kL=14.5$ and $\rho=0.44$ and its vertical cross-section are shown in (Figs.
-3ab).
-We note that the sample sizes are limited by our computational resources and the complexity of the unstructured mesh generation.
+In the second case, layers with various volume fractions $\rho$ have larger diameter to thickness ratios and numbers of particles up to 19000.
+A layer sample with $X_{R}=32$, $X_{L}=14.5$ and $\rho=0.44$ and its vertical cross-section are shown in Figs.
+2ab.
 For the DGTD simulations, we generate meshes with the particle systems surrounded by the perfectly matching layer (PML) to simulate an open boundary.
@@
 Therefore, we use samples with two types of the side boundaries: a randomly perturbed one and a smooth boundary.
-In the first case defects of the scale of $\sim\lambda$ that break such symmetry are introduced.
-In the second case a smooth boundary forms a dielectric discontinuity between the disordered medium and free space and, generally, one can expect whispering-gallery (WG) modes due to total internal reflection and cylindrical symmetry [48].
+In the first case, defects of the scale of the wavelength $\sim\lambda$, breaking such symmetry, are introduced (Fig.
+2b).
+In the second, case a smooth boundary forms a dielectric discontinuity between the disordered medium and free space and, generally, one can expect whispering-gallery (WG) modes due to total internal reflection and cylindrical symmetry [51].
 We consider a non-absorbing dielectric material and vary the refractive index $n$ in the range of $2.0\leq n\leq 3.0$.
-3 Results and discussion
 Figure 1: Propagation of a continuous, focused beam through a dense and a sparse layers of irregular particles.
@@
 (b) and (d), steady-state near-field intensity distributions $|E|^{2}$ in XY, YZ and back-XZ planes for both layers.
-A white dotted line indicates a layer boundary with the diameter $kR=22.1$ and the thickness $kL=19.5$.
 The incident beam is $E_{x}$-polarized.
-The size of particles is $kr=1.1$, and the refractive index is $n=3.0$.
-The sparse structure shows a classic multiple-scattering pattern while dense packing strongly limits transverse spreading of the beam.
-An animation of a short pulse propagation, illustrating the transport mechanism in the dense case, is provided in the Supplementary material A.1.
-Diffusive propagation of light is characterized by a transverse spread of energy and exponential energy decay in time as a result of multiple scattering.
+The size of particles is $X_{r}=1.1$, and the refractive index is $n=3.0$.
+Diffusive propagation of light in sparse media with $l^{*}>\lambda$ is characterized by the transverse spread of energy and an exponential decay of intensity of transmitted light in time as a result of multiple scattering and diffraction by individual scatterers.
 The geometry of the field spread experiences a change when particles are densely packed.
-This effect depends on the particle size and refractive index.
-Generally, at dense packing, far-field single scattering becomes impossible, and classic multiple scattering turns into percolation of light along random transmission channels [28, 29, 30].
-Depending on the particle size and degree of disorder, this can lead to formation of a Caylee-tree-like channel [28] or a transversely localized propagation [38, 36].
-This may occur even at a relatively small refractive index $n=1.5$ and $l^{*}\lesssim\lambda$ [28], which, however, should be insufficient for AL.
-Nevertheless, the diffuse field component appears to be significantly reduced under such conditions, which can be seen in the polarimetric measurements of backscattering from high-albedo powders with different packing densities [49, 50, 51].
-Little free space between the neighboring scatterers in a medium reduces the role of diffraction after each scattering event and, consequently, the contribution of the diffuse field [28].
-The measured polarimetric features produced by single particles on the surface, that are suppressed by unpolarized diffuse multiple scattering in sparse samples, get enhanced in compressed ones.
-Hence, in the search for real AL, we consider samples with parameters that are close to those predicted in the theoretical works.
+Generally, at dense packing, far-field scattering becomes impossible, and classic multiple scattering turns into percolation of light along random transmission channels [36, 37, 38].
+This can lead to formation of Caylee-tree-like channels [36] or a transversely localized propagation [23, 24], depending on the particle size and volume fraction.
+This transition may occur even at a relatively small refractive index $n=1.5$ and $l^{*}\lesssim\lambda$ [36], which, however, should be insufficient for AL.
 In Fig.
-1, we study focused beam propagation through a thick layer of particles with a size $kr=1.1$ and a refractive index $n=3.0$.
+1, we study focused beam propagation through thick layers of particles with a size $X_{r}=1.1$ and a refractive index $n=3.0$.
 The internal field patterns are compared for a dense and a sparse structures of $\rho=0.48$ (Fig.
@@
 In the sparse case (Fig.
-1d), we observe classic scattering by particle groups and individual particles, which appear to be weak scatterers at large separations from each other.
-In a dense system, the size of voids is reduced to $kl^{*}\lesssim 1$, and a focused beam tends to preserve its diameter along the path, re-emitting transmitted energy in the center of the back side.
+1d), we observe classic scattering by particle groups and individual particles.
+In a dense system, the size of voids is reduced to the scale of the wavelength or even smaller, and a focused beam tends to preserve its diameter along the path, re-emitting transmitted energy in the center of the back side.
 Small isotropic spreading occurs due to the leakage of energy through the evanescent field coupling.
 Therefore, the intensity of the beam decreases, but its geometric cross-section remains constant at the distance of the layer thickness.
-The dynamics of the process can be seen in the animation in the Supplementary material A.1.
-It shows the evolution of a focused beam, percolating through a dense layer without transverse spreading.
-Light propagation evolves mainly through evanescent coupling between individual scatterers without conventional free-space transport of the scattered waves.
+Excitation of the $E_{y}$ and, particularly, longitudinal $E_{z}$ field components (Fig.
+S3 in the Supplementary material III) proves that this kind of propagation is not equivalent to the wave transport in a homogeneous medium.
+The dynamics of the process can be seen in Video III of the Supplementary material III.
+It shows the evolution of a short pulse, percolating through a dense layer without transverse spreading.
+Light propagates mainly through evanescent coupling between individual scatterers without conventional free-space transport of the scattered waves.
 This reminds us of an evolution of a cellular automata (CA) in which the vector field state in an elementary cell at each time step is determined by its nearest environment, i.e.
 the shape of neighbour particles or the shape of a sub-wavelength void that is formed by the neighbors.
-The laws of interaction of the vector field and the air-dielectric interface become the CA rules.
+The laws of interaction of the vector field and the air-dielectric interface together with interference become the CA rules.
 Then, a sequence of such interactions determines the evolution of the entire vector field from the initial state, resulting in a forward pulse propagation.
-The fundamental property of the observed transport mechanism is that most of the energy of the propagating beam is concentrated in the free-space cavities, and little is refracted into the dielectric particles.
-A wave can couple to the neighbor-particle interface while being still coupled to the previous one propagating in a void between two interfaces.
-On the other hand, these interfaces are formally disconnected and belong to different scatterers.
-Thus, the field appears to be constantly coupled and one can say that there are no more ”scattering events”.
-Therefore, the term ”free path length” becomes irrelevant in its classic definition (see also [14]).
-$kl^{*}<1$ implies also an extremely small mean free time $t^{*}$, and the near field, being coupled to a particle, most often reaches its neighbour at time $t\ll T$, where $T$ is the wave oscillation period.
-Thus, we refer to the $kl^{*}$ parameter only to characterise the size of the voids in a particulate medium with respect to the wavelength.
-Furthermore, we note that the wave packet is partially preserved along the propagation path.
-Its effective wavelength becomes smaller, $\lambda_{eff}\approx 0.6\lambda_{inc}$, increasing the effective size of particles to $k_{eff}r\approx 1.8$.
-Nevertheless, the condition $kl^{*}\lesssim 1$ is preserved for most of the voids.
-Figure 2: Propagation of a continuous pulse through the dense layer shown in Fig.
-1a.
-Steady state distribution of the components of the electric near field $\vec{E}$ (XZ-plane) show that the transmitted pulse is not totally depolarized but, on the contrary, preserves most of its initial polarization state $E_{x}$ after transmission.
-The $E_{y}$ component as well as $E_{z}$ is excited locally in the voids between the particles along the propagation path but is not transmitted directionally.
-Diffusive propagation in a disordered medium depolarizes incident radiation after many scattering events that are strongly stochastic.
-A steady-state distribution of the electric field component $E_{x}$ in Fig.
-2 computed for a dense layer shows a counter-intuitive conservation of the original polarization state of the propagating beam at the exit.
-The $E_{y}$ and $E_{z}$ components are excited along the path and play a role in the field coupling process but they do not propagate directionally.
-This fact can explain the halt of transverse spreading: the propagating beam ”prefers” cavities where the initial polarization direction can be preserved and, hence, the $\vec{k}$-vector bending is avoided and the direction of propagation is not changed.
-Figure 3: (a) A layer of 19000 irregular particles packed with a volume fraction of 0.44 and with dimensions of $kR=32$ and $kL=14.5$.
-The size of particles is $kr=1.1$ and the refractive index is $n=3.0$.
+Figure 2: (a) A layer of 19000 irregular particles packed with a volume fraction of 0.44.
+The size of particles is $X_{r}=1.1$ and the refractive index is $n=3.0$.
 (b) vertical cross-section of the layer, showing intentional defects on the side boundary and the characteristic scale of voids $kl\lesssim 1$.
-(c) Normalized transmission of a short pulse by layers with volume fractions from $\rho=0.17$ to $0.5$ as a function of time $T(t)$, exhibiting a transition from the exponential regime to a non-exponential decay at longer times.
-(d) Time-dependent diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$.
+(c) Normalized transmission of a short pulse by layers with volume fractions from $\rho=0.17$ to $0.5$ as a function of time $T(t)$.
+(d) Diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$.
 Both axes have logarithmic scale.
-A dense layer with $\rho=0.44$ demonstrates a possible transition to $t^{-1}$ (dashed line) dependence at times $t>20t_{D}$.
-Figure 4: Simulation results for layers with different volume fractions and smooth side boundaries.
-The sample dimensions are of $kR=26$ and $kL=12$, respectively.
-The size of particles is $kr=1$ and the refractive index is $n=3.0$.
-(a) Normalized transmission of a short pulse by layers with volume fractions from $\rho=0.1$ to 0.5 as a function of time $T(t)$, exhibiting a transition from the purely exponential regime to a non-exponential decay at $\rho\gtrsim 0.3$.
-(b) Time-dependent diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$.
-Dense layers demonstrate a $t^{-1}$ (dashed lines) dependence at times $t>20t_{D}$.
-(c) and (d), Snapshots of the internal electric field (XY-cross-sections, $E_{x}$ component) distributions showing a whispering-gallery wave propagating along the boundary at time $t=25t_{D}$ for a dense ($\rho=0.44$) and a moderately sparse ($\rho=0.27$) layers.
-In the last case $T(t)$ is exponential despite the whispering-gallery effect.
-Transverse localization in a dense structure may be accompanied by the changes in the time dependence of transmission, analogous to those observed for dense packings of PEC spheres in [34].
+Dashed line shows a $t^{-1}$ dependence.
+Transverse localization in a dense structure may be accompanied by the changes in the time dependence of transmission, analogous to those observed for dense packings of metallic spheres in Ref.
+[22].
 Here we perform a related analysis.
@@
 the volume fraction and the refractive index.
-At first, we consider layers with perturbed side boundaries and with dimensions of $kR=32$ and $kL=14.5$.
+At first, we consider layers with perturbed side boundaries and with dimensions of $X_{R}=32$ and $X_{L}=14.5$.
 The volume fraction $\rho$ is varied between 0.17 and 0.5.
-[…]
+An example of a dense layer with $\rho=0.44$ and its cross-section are shown in Figures 2ab.
+AL is expected for materials with high refractive indices.
+E.g., TiO2 powders are reported as those where transverse localization and anomalous transmission can be observed [31].
+Therefore, we first fix the refractive index at $n=3.0$ to study the role of the volume fraction factor.
+The computed transmissions $T(t)$ for different samples are shown in Fig 2c.
+The layers are illuminated by a short Gaussian pulse.
+The arrival of the transmitted pulse intensity peak at the detector plane is recorded as time $t_{a}$ for each case.
+The horizontal axis represents normalized time $t/\penalty 50t_{a}$.
+The maximum observation time is limited to $\approx 25t_{a}$ as the samples are relatively thin and the measured intensity quickly reaches the numerical accuracy of double-precision numbers.
+However, we observe a non-exponential dependence of $T(t)$ for larger volume fractions near this limit.
+With increasing turbidity at larger $\rho$, we approach the $kl^{*}\lesssim 1$ condition obtaining not only a decreasing slope of $T(t)$ but a noticeable deviation of transmission from exponent at longer times.
 Non-exponential decay can be formally represented by means of a time-dependent diffusion coefficient $D(t)$.
-At the localized state, it decreases as $t^{-1}$ according to the solution of the diffusion equation for the case of $kl^{*}<1$ in time domain [39].
-Figure 3d shows time resolved diffusion coefficients $D(t)$ deduced from the computed $T(t)$ by local exponential fitting.
-One can see a transition from small variation of $D(t)$ in sparse structures to a steep function at $\rho>0.44$ and times $t>15t_{D}$.
+At the localized state, it decreases as $t^{-1}$ according to the solution of the diffusion equation for the case of $kl^{*}\lesssim 1$ in time domain [25, 22].
+Figure 2d shows time-resolved diffusion coefficients $D(t)$ deduced from the computed $T(t)$ by local exponential fitting.
+One can see a transition from small variation of $D(t)$ in sparse structures to a steep function at $\rho>0.44$ and times $t>15t_{a}$.
 Evidently we would have to extend the observation time limit in order to study the evolution of $T(t)$ at longer times and to find out if $D(t)$ can reach the predicted $t^{-1}$ dependence.
-This could be achieved by increasing sample thickness and diameter roughly by a factor of two or more.
-This would correspond to $kL\gtrsim 30$ and a number of particles $N>10^{5}$ which makes the problem computationally extremely expensive, even for modern large scale HPC clusters.
+This could be achieved by increasing sample dimensions roughly by a factor of two or more and, correspondingly the number of particles to $N\gtrsim 10^{5}$, which makes the problem computationally extremely expensive, even for modern large scale HPC clusters.
 Another way to hinder energy decay in time is to create conditions for a whispering-gallery effect at the side boundary of a cylindric layer.
-For this purpose we did computations for layers with smooth boundaries and a particle size of $kr=1$.
+For this purpose we did computations for layers with smooth side boundaries.
 The results for $T(t)$ and $D(t)$ dependencies at different volume fractions from $0.1$ to $0.5$ are shown in Figures 3ab.
-The dimensions of the layers are $kR=26$ and $kL=12$.
+The dimensions of the layers are $X_{R}=26$ and $X_{L}=12$.
 A densely packed sample with $\rho=0.44$ is shown in the inset in Figure 3a.
-Figures 4cd represent snapshots of the electric field distributions ($E_{x}$ field component) in the horizontal XY-plane at time $t=25t_{D}$ for a dense ($\rho=0.44$) and a moderately sparse ($\rho=0.27$) layers.
-The WG waves, excited by the incident pulse due to the smooth boundaries of samples with cylindric symmetry, are coupled to the side boundary due to total internal reflection and allow us reaching $t\approx 40t_{D}$ for relatively small samples.
+Figures 3cd represent snapshots of the electric field distributions ($E_{x}$ component) in the horizontal XY-plane at time $t=25t_{a}$ for a dense ($\rho=0.44$) and a moderately sparse ($\rho=0.27$) layers.
+The WG waves, excited by the incident pulse, allow us reaching $t\approx 40t_{a}$ for the samples with large volume fractions and relatively small thickness.
 At the same time the field distributions remain random in the bulk.
-Despite the WG effect, present in both cases, the structures show different transmission properties.
-The $T(t)$ plot in Figure 4a clearly shows a transition from purely exponential decay to localization at large volume fractions and times $t>20t_{D}$.
-This occurs at $\rho\gtrsim 0.3$.
-The WG effect is observed in all the layers above and below this value, even in the sparse structures for which the the condition $l^{*}\lesssim\lambda$ holds.
-This proves that it is not the WG mode itself but the particle packing density which is causing the non-exponential behaviour of $T(t)$.
-The diffusion coefficient $D(t)$ at times $t\gtrsim 20t_{D}$ obtained for the dense structures with $\rho\gtrsim 0.44$ follows a $t^{-1}$ fit in Figure 4b demonstrating that we have reached the AL regime.
-We note that this result is consistent, in terms of time and the volume fraction threshold, with that obtained for layers of PEC spheres in [34].
+Despite the WG effect, present in all cases, including the sparse structures, they show different transmission properties.
+The $T(t)$ plot in Figure 3a clearly demonstrates a transition from exponential decay to localization at large volume fractions $\rho\gtrsim 0.3$ and times $t>20t_{a}$.
+The diffusion coefficient $D(t)$ at times $t\gtrsim 20t_{a}$ obtained for the dense structures with $\rho\geq 0.44$ follows a $t^{-1}$ fit in Figure 3b demonstrating that we have reached the AL regime.
+This result is consistent, in terms of time and the volume fraction threshold, with that obtained for layers of metallic spheres in [22].
+Interestingly, a WG mode is established at times $t\approx 10t_{a}$, prior to the emergence of the $D(t)\sim t^{-1}$ regime.
+This proves that it is not the WG mode itself but the particle packing density which is causing the non-exponential behavior of $T(t)$.
+This is also confirmed by the results of simulations with periodic boundary conditions presented in the Supplementary material III.
+Figure 3: Simulation results for layers with different volume fractions and smooth side boundaries.
+The refractive index of the material is $n=3.0$.
+(a) Normalized transmission of a short pulse by layers with volume fractions from $\rho=0.1$ to 0.5 as a function of time $T(t)$.
+(b) Time-dependent diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$.
+Dense layers demonstrate a $t^{-1}$ (dashed lines) dependence at times $t>20t_{a}$.
+(c) and (d), Snapshots of the internal electric field (XY-cross-sections, $E_{x}$ component) distributions showing a whispering-gallery wave propagating along the boundary at time $t=25t_{a}$ for a moderately sparse ($\rho=0.27$) and a dense ($\rho=0.44$) layers.
+(e) Normalized spectra of transmitted light as functions of the dimensionless size parameter $X_{r}$ (inverse wavelength expressed relative to the particles size) for sparse $\rho=0.17$ and dense $\rho=0.44$ layers.
+Dashed line corresponds to a Lorentzian fit for $\rho=0.44$.
+The inset shows the Thouless conductance distribution $P(\textrm{ln}~g_{Th})$ for $\rho=0.44$ that has a mean value $\mu=-1.25$ and a standard deviation $\sigma=0.75$.
+To make the characterization of the localized transport more complete we compute transmission spectra for the sparse ($\rho=0.17$) and dense ($\rho=0.44$) samples (Figure 3e).
+For the horizontal axis a dimensionless size parameter $X_{r}$ is used, which is equivalent to the inverse wavelength relative to the particle size.
+The $X_{r}$ range in the plot spans approximately the full width at half maximum of the incident pulse.
+The $\rho=0.17$ spectrum is represented by broad, overlapping, maxima.
+The spectrum for $\rho=0.44$ consists of non-overlapping sharp peaks resulting from the states spatially confined in the medium [19, 26, 27, 28, 29].
+Dashed line shows a spectral fit obtained with an ensemble of Lorentzian oscillators.
+The separations of peaks are larger than their widths, which indicates that the Thouless criterion for AL is satisfied.
+With this, a Thouless conductance $g_{Th}$ can be defined as the ratio $g_{Th}=\delta\omega/\Delta\omega$, where $\delta\omega$ is the average width of two adjacent modes in the spectrum and $\Delta\omega$ is their spectral separation.
+Thus, $g_{Th}<1$ corresponds to a localized propagation.
+In the case of $\rho=0.44$, where the signatures of localization emerge, we estimate it as an average over all visible modes as $g_{Th}>\approx 0.36$.
+The inset in Figure 3e shows the distrubution $P(\textrm{ln}~g_{Th})$.
+It approximately follows a Gauss law according to the theoretical predictions [28].
 Figure 4: (a) Normalized transmission of a short pulse through a layer with a volume fraction $\rho=0.44$ and a smooth boundary as a function of time $T(t)$ at different refractive indices.
-(b) Time-dependent diffusion coefficient $D(t)$ as obtained by a local exponential fitting of $T(t)$.
-A localization signature $D(t)\sim t^{-1}$ emerges at times $t>20t_{D}$ and $n=3.0$.
-[…]
-Figure 6: A snapshot of the near field $|E|^{2}$ excited by the propagation of a short pulse through a dense layer with $\rho=0.48$ and $n=3.0$ shown in Fig.
-1a along $Z$ axis at times $t=2t_{D}$ (a) and $t=20t_{D}$ (b).
-A fraction of light is trapped by the disordered air-dielectric interface through evanescent coupling.
-This results in a constant flow along complex, highly disordered paths which slows down the energy decay through the layer boundaries.
-An animation showing the time dynamics of the near field can be found in the Supplementary material A.1.
-Our results imply a mechanism that slows down the energy decay at dense packing.
-Apparently, it should be connected with the percolation-like light transport in a thick, dense layer we analyzed above.
-Visualization of the field dynamics excited by a short pulse yields a clue to this mechanism.
-One can conventionally split up the propagating field into two components.
-The first one is the wave packet of the incident pulse that quickly diffuses through a layer.
-The second one is created by the energy leaked from the packet through the evanescent field-coupling.
-Being coupled once, this component remains trapped by the air-dielectric interface, which, together with the free-space cavities, represents a geometrically complex 3D structure.
-Despite little free space at dense packing, such a structure possesses a long-range connectivity, i.e.
-arbitrary points in a void can be connected with another point in an arbitrary free-space location.
-The paths connecting such arbitrary points can be much longer than the sample dimensions.
-In Fig.
-6, one can see snapshots of the near-field intensities inside the thick layer shown in Fig.
-1a at a time right after the pulse passage $t=2t_{D}$ and at the longer time $t=20t_{D}$.
-An animated field evolution for a period of time after $2t_{D}$ can be found in the Supplementary material A.1.
-The trapped field uniformly fills the sample volume and propagates along complex paths, constantly changing direction and forming hotspots in the voids.
-The paths can be not necessarily closed loops, but they are apparently sufficiently long.
-The field decay occurs only when the field reaches the sample boundaries.
-Thus, the energy in the interior needs substantial time in order to reach the boundary, much longer than the time of propagation along a straight line or a sequence of straight lines between scattering events in a sparse medium.
-The character of such an internal field evolution does not significantly change at longer times, and the decay of hotspots is not monotonous in each location.
-They demonstrate different behaviour in such a disordered system: those that are excited directly by the transmitted pulse decay while others arise in new locations.
-The decay rate is also different for different hotspots, i.e.
-they may have different life times.
-Some are decaying within $\approx 5$ oscillation periods while others are much more stable.
-Non-monotonous decay and enhancement with time may also take place in the same location, which proves a high geometrical complexity of the paths.
-If the energy of a path is leaked at the sample boundary, it is lost in the total decay.
-A leakage in the interior transfers energy to new paths with similar spatial scales and degrees of complexity, forming a complex percolation graph.
-Thus, the entire process appears to be a sophisticated spatiotemporal 3D flow of energy and should be a subject for further studies.
-In particular, such a mechanism can be relevant for low-absorbing materials as it involves field-interface interaction and propagation in free space.
-The role of the refractive index also becomes clear in this situation.
-It determines the balance between the energy transmitted by particles and chains of particles, enhancing a diffusive process, and the energy flowing through a random system of connected cavities with long path lengths that participates in localization.
-4 Conclusions
-We observe clear signatures of AL in the results of the full wave numerical simulations of light transmission by 3D dielectric particulate layers with high degrees of structural disorder.
-A transition from diffusive transport to localization occurs at approaching the Ioffe-Regel criterion for the mean free path length $kl^{*}\lesssim 1$ in densely packed systems and at the refractive index of $n=3.0$, which is consistent with the predictions of theoretical models.
-Simulation of the focused beam propagation in a dense layer demonstrates an absence of transverse spreading and anomalous conservation of the beam’s linear polarization state in contrast to the classical light scattering in a sparse analog structure.
-Time-resolved transmission of a short pulse in layers with similar densities and constituent particle sizes exhibits a non-exponential energy decay already at times $t>15t_{D}$.
-For observations at $t\gg t_{D}$ larger structures must be considered which makes the problem computationally very expensive even using HPC.
-Alternatively, we extend the observation time retaining sample cylindric symmetry.
-This allows excitation of whispering gallery modes that hinder field decay from a sample.
-Therewith we found that the time-resolved diffusion coefficient varies in agreement with the self-consistent theory of localization at times $t>20t_{D}$ and volume fractions $\rho\gtrsim 0.44$.
-Thus, Anderson localization of light is possible in dense highly disordered 3D dielectric structures which motivates a continued experimental search of the AL characteristics in uncorrelated disordered media, with realistic material refractive indices not exceeding $n=3.0$.
-[…]
-[…]
-[…]
-Methods We performed numerical simulations using a self-developed light scattering code based on the open source DGTD Maxwell solver MIDG published by Tim Warburton [52].
+(b) Diffusion coefficient $D(t)$ as obtained by a local exponential fitting of $T(t)$.
+Dashed line shows a $t^{-1}$ dependence.
+Another free parameter of the medium that is important for light transport is the refractive index of the material.
+To illustrate the transition between diffuse and localized regimes from this point of view, we do similar computations of $T(t)$ for the dense sample with $\rho=0.44$ and refractive indices ranging from n=n=2.0 to 3.0.
+Figure 4 shows that $n$ must be larger than 2.5 for a non-exponential decay, and further increase of $n$ enhances this effect at longer times.
+This is also consistent with theoretical estimations and experimental measurements for materials with different $n$ that suggest $n\gtrsim 2$ [9, 31].
+Thus, the refractive index contrast is another critical factor for AL in dielectric structures that can impede or enhance the localized transport.
+It turns out that $n$ should not be necessarily extremely large and a continued experimental search for AL in ’white paint’ powders like TiO2 is not futile.
+Our results suggest a mechanism that slows energy decay under dense packing conditions, likely linked to percolation-like transport.
+We provide our intepretation of the observed transition in the Supplementary material III , based on the analysis of the near-field dynamics.
+After a short pulse passes through the system, a decaying field component becomes trapped at the air–dielectric interface, exhibiting a complex three-dimensional spatiotemporal energy flow.
+In a strongly disordered medium with long-range connectivity, propagation paths can be intricate and significantly exceed the physical thickness of the sample.
+Coherent transport along such extended paths, emerging from an initially random near-field distribution, leads over time to the formation of stable, disorder-induced spatially localized modes.
+These modes have sufficiently long lifetimes and are responsible for the sharp peaks observed in the transmission spectrum.
+We note, that this highly complex field propagation cannot be adequately described by multiple light scattering in a cloud of identical point scatterers.
+In the last case, a vector treatment shows that near-field coupling and interference between the longitudinal and transverse components suppress AL [40, 39].
+However, it has also been shown that the amplitude of the longitudinal component, specific to point scatterers, non-monotonously depends on size for spheres and conditions can even be created to minimize it [52].
+In contrast to point-scatterers and spheres, irregular particles produce inherently non-symmetric dipolar near fields with random amplitudes determined by the particle orientation relative to the polarization of the incident wave.
+Consequently, the transverse and longitudinal components acquire random amplitudes, and their interference does not necessarily prevent AL.
+Nevertheless, the problem of the longitudinal field remains an important question worth a separate consideration.
+It would be particularly interesting to study its role in systems with different types of constituents and finely tunable spatial correlations.
+This can be very useful for understanding the dynamics of AL in 3D, especially since the local neighbourhood geometry is an important factor for an initial near field distribution and its subsequent evolution.
+In conclusion, our full-wave numerical simulations of light transmission through 3D disordered dielectric particulate layers reveal clear signatures of AL.
+We observe a transition from diffusive transport to localization as the system approaches the Ioffe–Regel condition at high volume fractions ($\rho\gtrsim 0.44$) and a refractive index of $n=3.0$, consistent with theoretical predictions.
+Simulation of focused beam propagation in a thick, dense layer demonstrates an absence of its transverse spreading.
+Time-resolved transmission of a short pulse in layers with similar volume fractions and particle sizes exhibits a non-exponential energy decay with the time-dependent diffusion coefficient varying in accordance with the self-consistent theory of localization.
+The transmission spectrum satisfies the Thouless criterion for AL in this case.
+Near-field visualizations reveal a clear dynamic transition from diffusive propagation to the formation of spatially localized modes.
+These findings demonstrate that Anderson localization of light can occur in highly disordered 3D dielectric structures, motivating further experimental efforts to identify its hallmarks in realistic disordered media with refractive indices not exceeding n = 3.0.
+Acknowledgements.
+The authors gratefully acknowledge the computing time granted by the Paderborn Center for Parallel Computing (PC2).
+Supplementary material
+I Numerical model
+We performed numerical simulations using a self-developed light scattering code based on the open source DGTD Maxwell solver MIDG published by Tim Warburton [53].
 The CST Microwave Studio software package based on the Time Domain Finite Integration Technique was used for a qualitative benchmarking the DGTD simulation results for simplified setups.
-In our code we implemented plane-wave and focused beam lights sources, boundary conditions and material description.
+In our code, we implemented plane-wave and focused beam lights sources, boundary conditions and material description.
 Here, we use continuous and Gaussian pulses.
@@
 All the DGTD simulations were done on the HPC cluster Noctua 2 of Paderborn University using up to 30 cluster nodes per simulation.
-Appendix A Supplementary material
-A.1 Near field dynamics
-Animated evolution of the near field in time helps to better represent the field dynamics in such a complex structure as a discrete random medium consisting of densely packed irregular particles.
-To support our analysis of the static field snapshots in Figure 1b we additionally simulate propagation of a short focused pulse in the same dense layer sample (Figure 1a) with the same model parameters.
-Animation A.1 represents an $XZ$ cross-section of the computational domain with the near-field intensity distribution and shows a cellular-automata-like behavior of the local field hotspots in the sub-wavelength voids between the closely neighboring particles.
-Figure A.1: Animated propagation of a short focused pulse in a dense layer shown in Figure 1a.
-Animation A.2 demonstrates dynamics of the near field that is trapped by the air-dielectric interface of the constituent particles after propagation of a short plane-wave pulse in a short period near $t=2t_{D}$.
-At high degree of disorder and in the absence of large voids the evanescent coupling mechanism slows down the energy decay from the system.
-Figure A.2: Animated dynamics of the near field in a dense layer shown in Figure 1a excited by a short plane wave pulse.
-A.2 Pulse transmission by layer samples with reduced dimensions
+II Short pulse transmission by layer samples with reduced dimensions
 Here we simulate short-pulse transmission by layers with reduced dimensions in order to study the transition to AL increasing the geometrical complexity through the layer thickness and the number of particles.
@@
 It is based on the Time Domain Finite Integration Technique, so, one can qualitatively benchmark our DGTD results with a different numerical method and a different electromagnetic solver.
-We take a cylindrically symmetric sample with $kR=26$, $kL=12$ and $\rho=0.44$ as a reference and consider reduced structures with thicknesses $kL=6$ and 3 and an elementary bi-layer structure.
+We take a cylindrically symmetric sample with $X_{R}=26$, $X_{L}=12$ and $\rho=0.44$ as a reference and consider reduced structures with the same bulk volime fraction and thicknesses $X_{L}=6$ and 3 and an elementary bi-layer structure.
 The refractive index is fixed at $n=3.0$.
-The results of the simulated transmission $T(t)$ for four layers are presented in Figure A.3a.
-A bi-layer shows purely exponential decay even though a WG mode is excited.
-The case of $kL=3$ is equivalent to a doubling the thickness of the bi-layer sample.
+The results of the simulated transmission $T(t)$ for four layers are presented in Figure S1a.
+A bi-layer shows exponential decay even though a WG mode is excited.
+The case of $X_{L}=3$ is equivalent to a doubling the thickness of the bi-layer sample.
 This appears to be enough to make the $T(t)$ curve non-exponential.
-In all four cases a WG mode is excited allowing us reaching the $25-40t_{D}$ region even with such a small thickness.
-Further thickness increasing leads to stronger deviations of exponent and faster reduction of the diffusion coefficient $D(t)$ (Figure A.3b).
-However, only with $kL=12$ we obtain a qualitative change at longer times and reach the $t^{-1}$ fit.
-Figure A.3: (a) Normalized transmission of a short pulse through layers with cylindric symmetry and different thickness $kL$ as a function of time $T(t)$.
-The diameter and volume fraction are $kR=26$ and $\rho=0.44$, correspondingly, and the refractive index is $n=3.0$.
+In all four cases a WG mode is excited allowing us reaching times $t>20t_{a}$.
+Further thickness increasing leads to stronger deviations from exponent and more rapid reduction of the diffusion coefficient $D(t)$ (Figure S1b).
+However, only with $X_{L}=12$ we obtain a qualitative change at longer times and reach the $t^{-1}$ fit.
+The whispering-gallery field pattern that we observe in all samples with cylindrical symmetry at longer times is, in fact, a complex 3D oscillation both along the round side boundary and in the vertical direction (along $Z$) between the upper and bottom sides of the cylinder.
+This is indeed a whispering-gallery process caused by cylindrical symmetry but not entirely a classic WG-mode that is well known for, e.g., solid micro-discs.
+There is also a constant flow of energy in the central part of the sample volume that interacts with the WG waves which is not typical for solid structures with circular symmetries.
+The random flow in the central part and in the bulk is a result of constant evanescent field coupling to a random network of dielectric interfaces feeded by the WG-mode.
+Such a near-field propagation process deserves a separate study as, to our knowledge, phenomena of this kind have not been described in literature yet.
+Figure S1: (a) Normalized transmission of a short pulse through layers with cylindric symmetry and different thickness $X_{L}$ as a function of time $T(t)$.
+The diameter and volume fraction are $X_{R}=26$ and $\rho=0.44$, correspondingly, and the refractive index is $n=3.0$.
 The energy decay from a simple double layer of particles is exponential while thicker layers demonstrate increasingly non-exponential transmission.
 (b) Diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$.
-$D(t)$ quickly evolves from a constant to a $t^{-1}$ fit (dashed line) with increasing thickness.
+$D(t)$ evolves from a constant to a $t^{-1}$ fit (dashed line) with increasing thickness.
 (c) Vertical cross-sections of the samples with different thickness.
-Figure A.4: Normalized transmission of a short pulse through a small layer of 2000 particles packed with volume fraction $\rho=0.5$ (inset) and with dimensions $kR=15$ and $kL=3$ as a function of time $T(t)$ at different refractive indices.
+Figure S2: Normalized transmission of a short pulse through a small layer of 2000 particles packed with volume fraction $\rho=0.5$ (inset) and with dimensions $X_{R}=15$ and $X_{L}=3$ as a function of time $T(t)$ at different refractive indices.
 Simulations are done with the CST Microwave Studio software package.
-For the computations with the CST FIT solver we take a cylindric sample with thickness $kL=3$ (a four-particle thick layer) and reduced diameter $kR=15$.
+For the computations with the CST FIT solver we take a cylindric sample with thickness $X_{L}=3$ (a four-particle thick layer) and reduced diameter $X_{R}=15$.
 Due to parallelization restrictions of CST MWS we need to simplify the target system.
-With the particle size $kr=1$ and the volume fraction $\rho=0.5$ the number of particles becomes relatively small in this case, N=2000.
+With the particle size $X_{r}=1$ and the volume fraction $\rho=0.5$ the number of particles becomes relatively small in this case, N=2000.
 With fixed sample parameters we vary the material refractive index $n$ from 2.0 to 3.0.
@@
 The electromagnetic problem is solved in time domain simulating propagation of the short plane-wave pulse, similarly to our DGTD method setup.
-The result in Figure A.4 is qualitatively similar to the $T(t)$ refractive index dependence in Figure 5a obtained for a much larger system with the DGTD method.
-Summarizing all our results we note that increasing three parameters: layer thickness, volume fraction and refractive index act in very similar ways causing transition from an exponential dependence of $T(t)$ to a non-exponential behavior with qualitative change at longer times.
+The result in Figure S2 is qualitatively similar to the $T(t)$ refractive index dependence in Figure 4a obtained for a larger system with the DGTD method.
+III Analysis of the near field dynamics
+In this section we analyze light transport in a thick layer with a large number of particles, shown in Figure 1a, in more detail and consider a problem of light transmission in layers with reduced dimensions and periodic boundary conditions, in order to interpret our observations of the localization signatures presented in Figure 3.
+At first, we visualize the excited near-field distribution in the thick sample and animate its evolution in time.
+We additionally simulate propagation of a short focused pulse in the same layer sample with the same model parameters.
+Video III represents an $XZ$-cross-section of the computational domain with the near-field intensity distribution.
+It shows a CA-like behavior of the field hotspots in the sub-wavelength voids between the closely neighboring particles.
+We note that the wave packet is partially preserved along the propagation path.
+Its effective wavelength becomes smaller, $\lambda_{eff}\approx 0.6\lambda_{inc}$, increasing the effective size of particles to $X_{r,eff}\approx 1.8$.
+Nevertheless, the condition $kl^{*}\lesssim 1$ or, at least, $kl^{*}\sim 1$ is preserved for most of the voids.
+{video}
+[…]
+Additional characterization can be obtained from the analysis of the polarization state of the near field.
+Generally, diffusive propagation in a disordered medium depolarizes incident radiation after many scattering events that are strongly stochastic in a disordered multi-particle system.
+A steady-state distribution of the electric field component $E_{x}$ in Fig.
+S3 computed for the same dense layer and same conditions as in Figure 1b shows a counter-intuitive conservation of the original polarization state of the propagating beam at the exit.
+The $E_{y}$ and $E_{z}$ components are excited along the path and play a role in the field coupling process but they do not propagate directionally.
+This fact can explain the halt of transverse spreading: the propagating beam ”prefers” cavities where the initial polarization direction can be preserved and, hence, the $\vec{k}$-vector bending is avoided and the direction of propagation is not changed.
+Figure S3: Steady-state distributions of the three components of the electric field at the propagation of a continuous focused beam through the dense layer with $\rho=0.48$ and $n=3.0$ shown in Figure 1a.
+From the focused short- and continuous-beam simulations we see that the fundamental property of the observed transport mechanism in densely packed layers is concentration of the transported energy in the free-space cavities, while less is refracted inside the dielectric particles.
+A wave can couple to the neighbor-particle interface while being still coupled to the previous one propagating in a void between two interfaces.
+On the other hand, these interfaces are disconnected and belong to different scatterers.
+Thus, the field appears to be constantly coupled and one can say that there are no more classic ”scattering events”.
+Therefore, the term ”free path length” becomes irrelevant in its classic definition.
+The condition $kl^{*}\lesssim 1$ implies also an extremely small mean free time $t^{*}$, and the near field, being coupled to a particle, most often reaches its neighbour at time $t\ll T$, where $T$ is the wave oscillation period.
+We note also, that the Ioffe-Regel criterion should be considered as qualitative for the localization problems [15].
+The parameter $kl^{*}$ can be measured only indirectly and the existing methods for such indirect estimations should be, therefore, validated for the case of the vector field propagating in a dense 3D medium.
+This is, however, beyond the scope of this paper.
+Thus, we use $kl^{*}$ just to characterise the size of the voids in a particulate medium with respect to the wavelength.
+We only roughly estimate it geometrically and then relate it to the Ioffe-Regel condition.
+The estimation follows from a simple consideration that at volume fractions $\rho=0.45–0.5$, which is close to the theoretical maximum for mono-disperse irregular constituents, the characteristic size of voids between particles cannot be larger than the particle size.
+The particle sizes in our simulations are $X_{r}=1–1.1$ or $X_{r,eff}\approx 1.8$ with respect to the effective wavelength at $\rho=0.44$ and $n=3.0$.
+Correspondingly, the characteristic free path length cannot be larger than that.
+For the further explanation of the light transport process at longer times, it is convenient to split up the propagating field into two components.
+The first one is the wave packet of the incident pulse that quickly passes through a layer.
+The second one is created by the energy leaked from the packet through the evanescent field-coupling.
+A field evolution video (Video III) can better represent the dynamics of the arrested component in a short period of time after $2t_{a}$.
+Randomization of the propagation direction, required for localization at longer times, is realized through coupling to the random air-dielectric interface.
+Being coupled once, the delayed component fills the sample volume and remains trapped by the interface, which represents a geometrically complex 3D structure.
+Despite little free space at dense packing, a particulate layer possesses a long-range connectivity, i.e., an arbitrary point in a void can be connected with another point in an arbitrary free-space location.
+Thus, field propagates along complex paths, constantly changing direction and forming randomly distributed hotspots in the voids (Figure S4, $t=2t_{a}$).
+{video}
+[…]
+Figure S4: Snapshots of the normalized near field $|E|^{2}$ excited by the propagation of a short pulse through a dense layer with $\rho=0.48$ and $n=3.0$ shown in Figure 1a along $Z$ axis at times $t=2t_{a}$ (a) and $t=20t_{a}$ (b).
+Figure S5: Cross-sections of the computational domain showing normalized near field $|E|^{2}$ in the sample with cylindric symmetry (Fig. 3a) and $\rho=0.44$ at time $t=25t_{a}$.
+The field decay occurs only when the field reaches the sample boundaries.
+Thus, the energy in the interior needs substantial time in order to reach the boundary, much longer than the time of propagation along a straight line or a sequence of straight lines between scattering events in a sparse medium.
+The decay of hotspots is not monotonic in each location.
+They demonstrate different behavior in such a disordered system already at earlier times (Video III): those that are excited directly by the transmitted pulse decay while others arise in new locations.
+The decay rate is also different for different hotspots, i.e.
+they may have different life times.
+Some are decaying within $\approx 5$ oscillation periods while others are much more stable.
+Non-monotonic decay and enhancement with time may also take place in the same location, which proves a high geometrical complexity of the paths.
+If the energy of a path is leaked at the sample boundary, it is lost in the total decay.
+A leakage in the interior transfers energy to new paths with similar spatial scales and degrees of complexity, forming a complex percolation graph.
+Evanescent coupling plays a crucial role here.
+Small dielectric particles act as resonant scatterers and the evanescent field tails from their surfaces overlap with neighboring particles.
+Such overlap is phase-sensitive and provides coherent pathways in a dense system.
+The character of such an internal field evolution does not significantly change in time but the hotspot distribution eventually becomes inhomogeneous.
+This becomes apparent in Figure S4 at $t=20t_{a}$ and it is even more clearly seen in the near-field inside a cylindric sample, for which we observe a non-exponential tail in transmission at longer times.
+Figure S5 shows $XY$-, XZXZ- and $YZ$-cross-sections of the computational domain with the near-field intensity for the $\rho=0.44$ sample at $t=25t_{a}$.
+Together with the WG pattern along the boundary, one can see notice clustering of the intensity hotspots in the bulk.
+To make sure that such a structuring is preserved in the absence of the WG mode, we did a similar series of simulations for a square layer with periodic boundary conditions.
+All model parameters, such as gaussian pulse, particle size and refractive index ($n=3.0$) remain the same, except reduced dimensions of a unit cell ($X_{d}=18$ along X and Y axes and $X_{L}=9.6$ along Z).
+We vary volume fraction $\rho$ from 0.18 to 0.44.
+Periodic boundary conditions is another way to hinder energy decay from a finite sample.
+However, one should be careful and avoid phenomena like standing waves and guiding modes, that may appear over long time in an artificially infinite thin layer and influence the $T(t)$ curve.
+Figure S6: a.
+Normalized transmission of a short pulse through periodic layers with different volume fractions (inset: $\rho=0.44$) as a function of time $T(t)$.
+b.
+Diffusion coefficient $D(t)$ deduced from the $T(t)$ data.
+c.
+Transmission spectra for different volume fractions measured at longer times and at the normalized transmission level of $T\approx 10^{-9}$.
+d.
+XYXY-snapshots of the near-field intensity $|E|^{2}$ inside layers with different volume fractions at times corresponding to the normalized transmission $T\approx 10^{-8}$.
+{video}
+[…]
+First of all, we need to show that such a system reveals a non-exponential tail in $T(t)$ and a time-dependent diffusion coefficient $D(t)$ at longer times.
+Despite the small dimensions of the unit cell and small number of particles ($N=5000$ for $\rho=0.44$) we reach long enough observation times and see a qualitative transition in light transmission characteristics, very similar to that observed for larger finite cylindric samples.
+Figure S6 shows an emergence of a non-exponential tail in $T(t)$ at larger volume fraction ($\rho=0.44$) (a) and a time-dependent diffusion coefficient $D(t)$ (b).
+Transmission spectra for different volume fractions also show a transition from a number of overlapping peaks to isolated very sharp ones (Figure S6c).
+The horizontal axis in the spectral plots is scaled in the units of the particle size parameter, so that the pulse central wavelength corresponds to $X_{r}\approx 1.15$.
+Such a property of the dense system is sample-independent.
+We additionally simulated transmission for another realization of disorder for $\rho=0.44$ and it demonstrates qualitatively similar $T(t)$ dependence and transmission spectrum (Fig S7).
+Figure S7: Normalized transmission $T(t)$ (a) and transmission spectra (b) for two realizations of disordered layers with volume fraction $\rho=0.44$.
+Figure S8: Examples of the near-field intensity distributions $|E|^{2}$ at different time steps, shortly after $t=4t_{a}$ (a) and at $t>75t_{a}$ (b), when the transmission dependence $T(t)$ in Figure S6a becomes significantly non-exponential in the case of volume fraction $\rho=0.44$.
+Figure S6d shows distributions of the near-field intensity in $XY$-plane for different volume fractions recorded at different times, when normalized transmission reaches $T(t)\approx 10^{-8}$.
+Here one can also notice a qualitative change.
+We see dipolar excitations of single particles and multiple wave scattering in a sparse structure, whereas, in a dense packing, the field energy is concentrated in small voids.
+Moreover, one can see clusters of hotspots and dark regions for the $\rho=0.44$ sample.
+Now we can focus on the dense sample and track the changes in the near-field dynamics in time.
+Figure S8 represents field snapshots taken for the $\rho=0.44$ case at different time-steps before and after the $T(t)$ curve becomes non-exponential.
+The upper raw (Figure S8a) corresponds to the times shortly after $t=4t_{a}$ and shows randomly scattered hotspots, indicating diffusion.
+$XY$ and $XZ$ field maps in Figure S8b are taken at $t>75t_{a}$ when $T(t)$ is non-exponential and $D(t)$ is significantly time-dependent.
+At such long times field energy distribution turnes into a structured pattern of interacting clusters of hotspots.
+These clusters are, in fact, resonances induced by coherent propagation along highly disordered pathways.
+Video III shows their dynamics.
+The clusters (as well as dark regions) evolve in time but remain stable for tens of oscillation periods which is consistent with the sharp-peak spectrum.
+Single hotspots may have lifetimes of 10-20 periods.
+The lifetimes of the modes in the spectrum in Figure S6d ($\rho=0.44$) can be roughly estimated from the peak widths as long as $\sim 100$ periods which is a strong indication of localization.
+To quantify this dynamics we calculated the inverse participation ratio (IPR) for such field maps for different volume fractions and different times using the $XY$-cross-section data.
+For a 2D scalar field $z(x,y)$ it is typically defined as
+$\textrm{IPR}=\frac{\sum_{i}{z_{i}}^{2}}{{\left(\sum_{i}{z_{i}}\right)}^{2}}.$ (1)
+$z$ is the near-field intensity in our case and IPR gives a measure of localization: lower IPR means energy spread out over many points and higher IPR implies localization at few points.
+Figure S9: Time depence of the inverse participation ratio calculated from the intensity maps for different volume fractions.
+Linear regression indicates increasing energy concentration in loclaized modes with time.
+Dashed lines show a 95% confidence interval.
+The calculated IPR as a function of time shows persistent growth for the volume fraction $\rho=0.44$ in Figure S9, that confirms our observation.
+Interestingly, this measure reflects more uniform spread of energy in sparse samples for the same measurement period.
+At $\rho=0.44$, field becomes increasingly concentrated in smaller spatial regions and the system evolves towards stronger spatial localization over time.
+The contrast between bright clusters of concentrated field hotspots and dark regions persistently grows.
+Therefore, such a process is not stationary and the observed resonances are not eigenmodes of the system but rather a time-dependent superposition of multiple localized modes.
+The excited modes can be initially isolated but their superposition and interference between them creates apparent spatial connections.
+Thus, we can conclude that we observe, for the first time, a real-time dynamics of Anderson localization formation in 3 dimensions.
+We can give now a general interpretation of our results.
+Initial excitation by a short pulse creates superposition of many modes.
+Light explores many pathways, some of them begin to dominate due to interference while others decay.
+Obviously, extended modes decay faster then localized or quasi-localized ones.
+When the total energy of localized modes is significant enough, hotspot clusters become distinct and we obtain a slowdown in $T(t)$.
+After that, the system evolves towards pure Anderson modes with infinite lifetimes, which cannot be eventually reached due to the finiteness of the sample and its small thickness.
+The localization length can be roughly estimated from the field images at different times as a few particle lengths or $\xi\approx 1-2\lambda$.
+This is also consistent with the layer thickness test in Figure S1.
+Individual modes can localize at the layer thickness $X_{L}=9.6$ ($L>\xi$) but the system is not large enough for efficient mode isolation ($L$ is not $\gg\xi$).
+Most of them extend across a significant fraction of the layer thickness which leads to spatial overlaps and interactions with the layer boundaries, increasing energy leakage.
+In fact, our simulation shows how Anderson localization manifests itself in a finite system.
+In order to directly observe a high-quality mode isolation we need a layer thickness of at least $L\sim 10\xi$ or $L\sim 10-20\lambda$ and with $\sim 10^{5}$ particles.
+This would require very large computing power, which is achievable, in principle, in large-scale massively parallel simulations.
+With our largest samples we get closer to this regime as it can be seen in Figure S1.
+Thus, it is not enough just to reach long observation times like we do this incorporating the WG-effect or periodic boundary conditions.
+Providing sufficiently large 3D volume for isolation of the disorder-induced localized modes is important.
+We note, that the mechanism described above can be relevant also for low-absorbing materials as it involves mostly field-interface interaction and propagation in free space.
+The role of the real part of the complex refractive index becomes clear in this situation.
+It determines the balance between the energy transmitted by particles and groups of particles, enhancing a diffusive process, and the energy flowing through the random network of connected voids with long path lengths that enforces localization.